need help evaluating x=-4 , y=6, z=-8
2x+3y
5xy-z
-z+2x
_______________________________
p-12= -15
-11= -4+d
-1.6k= -15
not from any textbook that I know of, from a 10 page worksheet

Answer by midwood_trail(310)

Evaluate:
x = -4 , y = 6, z = -8
2x+3y
5xy-z
-z+2x
Replace x, y and z with their given values in each expression and simplify.
I will do one of them.
2x + 3y
Let x = -4 and Let y = 6
Then 2(-4) + 3(6) = -8 + 18 = 10
Can you do the rest now?
_______________________________

p-12= -15
-11= -4+d
-1.6k= -15
Here, you need to solve each equation for the given variable.
I will do one of them.
p - 12 = -15
Add 12 to both sides.
p = 12 - 15
p = -3
Can you do the rest now?
I received your e-mail.
Here is the full reply:
Evaluate using: x = -4 , y = 6, z = -8


2x+3y...Replace x with -4 and y with 6


2(-4) + 3(6) = -8 + 18 = 10


==================


5xy-z...5(-4) times (6) -(-8) = 5(-24) + 8 = -120 + 8 = -112


==================



-z+2x...-(-8) + 2(-4) = 8 - 8 = 0


Did you follow?

_______________________________

When crossing the equal sign, change the sign infront of the numbers to the opposite ONLY when adding and subtracting.


p -12 = -15...We add 12 to both sides.


p = +12 - 15


p = -3


========================

-11= -4 + d...Here, we add 4 to both sides.


-11 + 4 = d


-7 = d


====================

-1.6k= -15....Here, we apply the division property or rule of algebra. To find k, divide both sides by the coefficient of k, which is the decimal number -1.6 in this question.


So, k = -15 divided by -1.6


Then, k = 9.375


Did you follow?